# Performance Modeling for Correlation-based Neural Decoding of Auditory Attention to Speech

### v1

@@ Abstract
To address this, we propose a new method to model the complete performance curve of correlation-based AAD algorithms using (known or measured) correlation statistics for only one decision window length. This enables continuous and efficient monitoring of the trade-off between accuracy and decision window length, without having to evaluate the model for all possible window lengths. Furthermore, the proposed technique provides a theoretical model to inform new correlation-based AAD algorithm design by translating target accuracy requirements into corresponding correlation coefficient targets, and vice versa. As this is an accuracy evaluation framework, we assume ground-truth attention labels are known, and leave the integration with unsupervised accuracy estimation as in [6] for future work.
@@ II-B Modeling of correlations in the AAD decision system
In this paper, we evaluate the AAD accuracy assuming the ground-truth labels are known, as opposed to the unsupervised accuracy estimation technique in [6]. For a given window length and sampling frequency , the measured correlation coefficient over samples between the (zero-mean) decoded response and encoded attended speech signal corresponds to:
@@
Since the AAD decision rule compares per-window attended and unattended correlations ( and ), and given that the inverse hyperbolic tangent is monotonically increasing, we can use instead of in Fig. 1 to make the binary decision on attention. Under the assumption that and are uncorrelated (as empirically validated in [6]), we can use as our decision variable. The AAD accuracy can then be computed using the cumulative distribution function (CDF) of the normal distribution, evaluated in zero [6]:
@@ III-A Linear decoding on Das-2016 dataset
This AAD algorithm is evaluated on the publicly available Das-2016 dataset, which contains EEG recordings from 16 participants attending to one of two competing speakers for [14]. We compute the attended and unattended correlations using user-specific 5-fold cross-validation on trials. Ground-truth AAD accuracies and (un)attended correlations are computed across decision window lengths of s. To analyze the impact of the amount of estimation data, we vary the amount of data used ( min) by randomly sampling from the available test correlations, which is repeated 10 times per condition. This AAD algorithm and evaluation serves as the default experimental setup in Section IV unless specified otherwise.
@@ IV-A Average performance curve estimation
The first overall observation is that the modeled performance curves generally closely approximate the ground truth. As expected, optimal prediction accuracy is observed at the corresponding baseline window length, confirming the modeling of the AAD decision variable using a normal distribution after the Fisher transformation. Prediction error increases with the extrapolation distance between baseline and target window lengths. The baseline window length can, therefore, be tailored towards the range of window lengths of interest (for example, the shorter ones in the context of AAD gain control systems [4]). Generically, a good choice is in the middle of the considered range of window lengths. Here, based on Fig. 2, we select as the baseline for subsequent analyses.
@@ IV-C Individual performance curve estimation
The results clearly show that the accuracy prediction method closely approximates the ground-truth accuracy, even on a per-participant level, with a mean absolute error of pp (standard deviation ). The estimated -confidence intervals contain the true accuracy in of cases, validating our uncertainty quantification.
@@ IV-D Validation on VLAAI network on Fuglsang-2018 dataset
We validate the performance modeling method on the second experimental setup with the VLAAI deep neural network on the Fuglsang-2018 dataset (Section III-B). Informed by the previous results, we use a baseline decision window length and of estimation data, again with 10 repetitions.
Similarly to Fig. 2, Fig. 4 shows the modeled average performance curve across participants vs. the ground-truth one. Table II shows the per-participant mean and standard deviation of the estimation errors across target decision window lengths and repetitions. Both demonstrate comparable performance to the first setup on the Das-2016 dataset, with a mean absolute error across participants of pp (standard deviation ), and coverage by the estimated -confidence interval.

### v2

@@ Abstract
To address this, we propose a new method to model the complete performance curve of correlation-based AAD algorithms using (known or measured) correlation statistics for only one decision window length. This enables continuous and efficient monitoring of the trade-off between accuracy and decision window length, without having to evaluate the model for all possible window lengths. Furthermore, the proposed technique provides a theoretical model to inform new correlation-based AAD algorithm design by translating target accuracy requirements into corresponding correlation coefficient targets, and vice versa. As this is an accuracy evaluation framework, we assume ground-truth attention labels are known, and leave the integration with unsupervised accuracy estimation as in López-Gordo et al. [6] for future work.
@@ II-B Modeling of correlations in the AAD decision system
In this paper, we evaluate the AAD accuracy assuming the ground-truth labels are known, as opposed to the unsupervised accuracy estimation technique in López-Gordo et al. [6]. For a given window length and sampling frequency , the measured correlation coefficient over samples between the (zero-mean) decoded response and encoded attended speech signal corresponds to:
@@
Since the AAD decision rule compares per-window attended and unattended correlations ( and ), and given that the inverse hyperbolic tangent is monotonically increasing, we can use instead of in Fig. 1 to make the binary decision on attention. Under the assumption that and are uncorrelated (as empirically validated in López-Gordo et al. [6]), we can use as our decision variable. The AAD accuracy can then be computed using the cumulative distribution function (CDF) of the normal distribution, evaluated in zero [6]:
@@ III-A Linear decoding on Das-2016 dataset
This AAD algorithm is evaluated on the publicly available Das-2016 dataset, which contains EEG recordings from 16 participants attending to one of two competing speakers for [14]. We compute the attended and unattended correlations using user-specific 5-fold cross-validation on trials. Ground-truth AAD accuracies and (un)attended correlations are computed across decision window lengths of s. To analyze the impact of the amount of estimation data, we vary the amount of data used ( min) by randomly sampling from the available test correlations, which is repeated ten times per condition. This AAD algorithm and evaluation serves as the default experimental setup in Section IV unless specified otherwise.
@@ IV-A Average performance curve estimation
The first overall observation is that the modeled performance curves generally closely approximate the ground truth. As expected, optimal prediction accuracy is observed at the corresponding baseline window length, confirming the modeling of the AAD decision variable using a normal distribution after the Fisher transformation. Prediction error increases with the extrapolation distance between baseline and target window lengths. The baseline window length can, therefore, be tailored towards the range of window lengths of interest (for example, the shorter ones in the context of AAD gain control systems [4]). In general, a good choice is in the middle of the considered range of window lengths. Here, based on Fig. 2, we select as the baseline for subsequent analyses.
@@ IV-C Individual performance curve estimation
The results clearly show that the accuracy prediction method closely approximates the ground-truth accuracy, even on a per-participant level, with a mean absolute error of pp (standard deviation pp). The estimated -confidence intervals contain the true accuracy in of cases, validating our uncertainty quantification.
@@ IV-D Validation on VLAAI network on Fuglsang-2018 dataset
We validate the performance modeling method on the second experimental setup with the VLAAI deep neural network on the Fuglsang-2018 dataset (Section III-B). Informed by the previous results, we use a baseline decision window length (closest to in the range of window lengths in Section III-B) and of estimation data, again with ten repetitions.
Similarly to Fig. 2, Fig. 4 shows the modeled average performance curve across participants vs. the ground-truth one. Table II shows the per-participant mean and standard deviation of the estimation errors across target decision window lengths and repetitions. Both demonstrate comparable performance to the first setup on the Das-2016 dataset, with a mean absolute error across participants of pp (standard deviation pp), and coverage by the estimated -confidence interval.